OpenStudy (anonymous):

find the coefficient of x^3 in the expansion of (3x+2)^8

OpenStudy (anonymous):

\[\dbinom{8}{3}(3x)^32^5\] is the term you are looking for

OpenStudy (anonymous):

\[\dbinom{8}{3}=\frac{8\times 7\times 6}{3\times 2}=8\times 7=56\] \[3^3=27\] \[2^5=32\] multiply those numbers together to get your coefficient

OpenStudy (anonymous):

how do you find the first part ?

OpenStudy (anonymous):

Since the power 8, and the third element is the one you pick on!

OpenStudy (anonymous):

i don't know what "the first part" is

OpenStudy (anonymous):

if you mean \(\binom{8}{3}\) that is because you want \(x\) to the power of 3, and you are raising this binomial to the power of 8